Two boilers one with super heater and other without super heater are delivering equal quantities of steam into a common main. The pressure in the boilers and main is18 bar. The temperature of steam from a boiler with a super heater is 350°C and temperature of

the steam in the main is 250°C.

Determine the quality of steam supplied by the other boiler. Take Cpsteam = 2.25 kJ/kgK.


1
Expert's answer
2022-01-24T15:39:02-0500



Let the quality of steam supplied by the other boiler be x

Moisture in 1kg "=(1-x)" kg



From saturated steam table at 18 bar;

(a) Saturated steam temperature "=209.8\u00b0C"

(b) Latent heat of steam "(h_{fg})=1900.28kJ\/kg"



Heat gained by wet steam "=" Heat lost by superheated steam


"(1-x)h_{fg}+c_{ps}(t_m-t_s)=c_{ps}(t_{sup}-t_m)"


"(1-x)1900.28+2.25(250-209.8)=2.25(350-250)"


Solving

"x=0.9219"